# Hodge structures on conformal blocks

### v1

@@ 7.2. Modular functors are defined over number fields
In this subsection, we deduce from the well known fact that all complex ribbon categories can be defined over number fields that any complex modular functor $\mathcal{V}$, whose associated weakly ribbon category $\mathrm{C}_{\mathcal{V}}$ is ribbon, can be defined over a number field.
@@
Any ribbon category over $\mathbb{C}$ is defined on a number field.
@@
Let $\mathcal{C}=(\mathrm{Vect}_{\Lambda},\otimes,c,\beta,\theta)$ be a ribbon category over $\mathbb{C}$. By the theorem, the underlying fusion category is defined over a number field $K$. So $c$ is defined over $K$. Now, the data of the braiding $\beta$ is equivalent to the data of a certain functor of multi-fusion categories $\Phi:\mathcal{C}\times\mathcal{C}\rightarrow\mathcal{C}$ called a multiplication (see Yetter’s article [24, 2.18,2.19] for a definition and details). By the theorem, $\Phi$ is defined over a finite extension $K^{\prime}$ of $K$. Hence so is $\beta$. The data of the twist $\theta$ is the same as that of the numbers $t_{\lambda}\in\mathbb{C}^{*}$. By the Anderson-Moore-Vafa theorem, the twists $t_{\lambda}$ are always roots of unity, and thus defined on a finite extension $K^{\prime\prime}$ of $K^{\prime}$ (see, for example [12, 8.18.2]). In conclusion, $\mathcal{C}$ is defined over the number field $K^{\prime\prime}$. ∎
@@
Any modular, genus 00 modular or rooted functor $\mathcal{V}$ over $\mathbb{C}$ whose weakly ribbon category $\mathrm{C}_{\mathcal{V}}$ is ribbon is defined over a number field.
@@
In our context, this means that there exists an isomorphism $\varphi:i_{1}^{*}\mathrm{C}_{\mathcal{V}}\simeq i_{2}^{*}\mathrm{C}_{\mathcal{V}}$ of ribbon categories over $\mathbb{C}\otimes_{K}\mathbb{C}$ satisfying the cocycle conditions. If $\mathcal{V}$ is modular, its central charge $c\in\mathbb{C}^{*}$ is a root of unity (2.37). Hence we may assume that $c$ is in $K$. By Theorem 5.6, $\varphi$ extends to an isomorphism $\widetilde{\varphi}:i_{1}^{*}\mathcal{V}\simeq i_{2}^{*}\mathcal{V}$ of modular, genus 00 modular or rooted functors over $\mathbb{C}\otimes_{K}\mathbb{C}$ satisfying the cocycle conditions. By descent, this means that $\mathcal{V}$ has a $K$-lattice. ∎
7.3. Proof of existence and uniqueness of Hodge structures on modular functors
7.3.1. Proof of Theorem 3.21
Let $\mathcal{V}$ be a complex modular functor with semisimple quantum representations such that $\mathrm{C}_{\mathcal{V}}=(\mathrm{Vect}_{\Lambda},\otimes,c,\beta,\theta)$ is ribbon. Then by the non-abelian Hodge correspondence for complex modular functors (6.15), we have a modular functor of Higgs bundles that we denote $(\mathcal{V},\theta)$.
For $t\in\mathbb{R}_{+}^{*}$, $(\mathcal{V},t\theta)$ is also a modular functor of Higgs bundles and by the correspondence we have a family of modular functors $(\mathcal{V}_{t})_{t\in\mathbb{R}_{+}^{*}}$. By Proposition 6.13, this family, seen as a family of gluing and vacuum isomorphisms, is continuous (for the product topology).
Notice that ¯$\overline{\mathcal{M}}_{0}^{2}(r)\simeq\mathrm{B}\mu_{r}$ (see our convention 2.25) and ¯$\overline{\mathcal{M}}_{0}^{3}(r)\simeq\mathrm{B}\mu_{r}^{3}$, so that $\theta=0$ on these spaces. Hence for any $t$, since $(\mathcal{V},\theta)$ and $(\mathcal{V},t\theta)$ coincide on these spaces so do $\mathcal{V}$ and $\mathcal{V}_{t}$. In particular, the associated weakly ribbon categories $\mathrm{C}_{\mathcal{V}}$ and $\mathrm{C}_{\mathcal{V}_{t}}$ have the same braiding and twist, and may differ only in their associator, that we denote $c$ and $c_{t}$ respectively.
Being ribbon is an open algebraic condition on the associator $c$, see for example [4, 5.3.13]. Hence, as $\mathrm{C}_{\mathcal{V}}$ is ribbon, there exists $\epsilon>0$ such that $\mathrm{C}_{\mathcal{V}_{t}}$ is ribbon for $t\in(1-\epsilon,1+\epsilon)$.
By Corollary 7.7, $(c_{t})_{t\in(1-\epsilon,1+\epsilon)}$ is a trivial deformation of $c=c_{1}$ and there exists a continuous family of isomorphisms $f_{t}:\mathrm{C}_{\mathcal{V}}\rightarrow\mathrm{C}_{\mathcal{V}_{t}}$ with $f_{1}=\mathrm{id}$. By fullness in Theorem 5.6, these isomorphisms lift to isomorphisms $\mathcal{V}\rightarrow\mathcal{V}_{t}$. Then, by the non-abelian Hodge correspondence for complex modular functors (6.15), these isomorphisms induce isomorphisms $f_{t}:(\mathcal{V},\theta)\rightarrow(\mathcal{V},t\theta)$.
@@ 7.3. Proof of existence and uniqueness of Hodge structures on modular functors
By the correspondence (6.15) they induce automorphisms $k_{t}$ of $\mathcal{V}$ and thus also of the ribbon category $\mathrm{C}_{\mathcal{V}}$. By Corollary 7.10, there exists a family $\delta_{t}:\Lambda\rightarrow\mathbb{C}^{*}$ with $\delta_{t}(0)=0$ such that $k_{t}$ is on $\hom([\lambda],[\mu]\otimes[\nu])$ multiplication by $\delta_{t}(\mu)\delta_{t}(\nu)\delta_{t}(\lambda)^{-1}$. Thus $k_{t}$ is on $\mathcal{V}_{g}(\underline{\lambda})$ multiplication by $\prod_{i}\delta_{t}(\lambda_{i})$.
@@
7.3.2. Proof of Corollary 3.22
The proof essentially follows from that of [23, Theorem 5]. See there for details. Let $\mathcal{V}$ be a complex modular, genus 00 modular or rooted functor such that $\mathrm{C}_{\mathcal{V}}$ is ribbon. Denote by $\Lambda$ its set of colors. Fix $g,n\geq 0$ tangent stable and $\underline{\lambda}\in\Lambda^{n}$. Denote by $\rho_{g}(\underline{\lambda})$ the monodromy representation of the local system $\mathcal{V}_{g}(\underline{\lambda})$, and by $\Gamma$ the fundamental group of ¯$\overline{\mathcal{M}}_{g}^{n}(r,s)$.

### v2

@@ 7.2. Modular functors are defined over number fields
In this subsection, we deduce from the well known fact that all complex ribbon fusion categories can be defined over number fields that any complex modular functor $\mathcal{V}$ can be defined over a number field.
@@
Any braided or ribbon fusion category over $\mathbb{C}$ is defined on a number field.
@@
Let $\mathcal{C}=(\mathrm{Vect}_{\Lambda},\otimes,c,\beta)$ be a braided fusion category over $\mathbb{C}$. By the theorem, the underlying fusion category is defined over a number field $K$. So $c$ is defined over $K$. Now, the data of the braiding $\beta$ is equivalent to the data of a certain functor of multi-fusion categories $\Phi:\mathcal{C}\times\mathcal{C}\rightarrow\mathcal{C}$ called a multiplication (see Yetter’s article [31, 2.18,2.19] for a definition and details). By the theorem, $\Phi$ is defined over a finite extension $K^{\prime}$ of $K$. Hence so is $\beta$.
If $\mathcal{C}$ is ribbon for a twist $\theta$, the data of such a twist is the same as that of the numbers $t_{\lambda}\in\mathbb{C}^{*}$. By the Anderson-Moore-Vafa theorem, they are always roots of unity, and thus defined on a finite extension $K^{\prime\prime}$ of $K^{\prime}$ (see, for example [12, 8.18.2]). In conclusion, $\mathcal{C}$ is defined over the number field $K^{\prime\prime}$. ∎
@@
Any modular, genus 00 modular, ribbon or braided functor $\mathcal{V}$ over $\mathbb{C}$ is defined over a number field.
@@
In our context, this means that there exists an isomorphism $\varphi:i_{1}^{*}\mathrm{C}_{\mathcal{V}}\simeq i_{2}^{*}\mathrm{C}_{\mathcal{V}}$ of ribbon categories over $\mathbb{C}\otimes_{K}\mathbb{C}$ satisfying the cocycle conditions. If $\mathcal{V}$ is modular, its central charge $c\in\mathbb{C}^{*}$ is a root of unity (2.40). Hence we may assume that $c$ is in $K$. By Theorem 5.6, $\varphi$ extends to an isomorphism $\widetilde{\varphi}:i_{1}^{*}\mathcal{V}\simeq i_{2}^{*}\mathcal{V}$ of modular, genus 00 modular, ribbon or braided functors over $\mathbb{C}\otimes_{K}\mathbb{C}$ satisfying the cocycle conditions. By descent, this means that $\mathcal{V}$ has a $K$-lattice. ∎
8. Proof of existence and uniqueness of Hodge structures on modular functors
8.1. Proof of Theorem 3.19
Let $\mathcal{V}$ be a complex modular, genus 00 modular, ribbon or braided functor. Then by the non-Abelian Hodge correspondence for complex modular functors (6.15), we have a functor of Higgs bundles that we denote $(\mathcal{V},\theta)$.
For $t\in\mathbb{R}_{+}^{*}$, $(\mathcal{V},t\theta)$ is also a functor of Higgs bundles and by the correspondence we have a family of functors $(\mathcal{V}_{t})_{t\in\mathbb{R}_{+}^{*}}$. By Proposition 6.13, this family, seen as a family of gluing and vacuum isomorphisms, is continuous (for the product topology).
Notice that ¯$\overline{\mathcal{M}}_{0}^{2}(r)\simeq\mathrm{B}\mu_{r}$ (see our convention 2.26) and ¯$\overline{\mathcal{M}}_{0}^{3}(r)\simeq\mathrm{B}\mu_{r}^{3}$, so that $\theta=0$ on these spaces. Hence for any $t$, since $(\mathcal{V},\theta)$ and $(\mathcal{V},t\theta)$ coincide on these spaces so do $\mathcal{V}$ and $\mathcal{V}_{t}$. In particular, the associated ribbon or braided fusion categories $\mathrm{C}_{\mathcal{V}}$ and $\mathrm{C}_{\mathcal{V}_{t}}$ have the same braiding (and twist), and may differ only in their associator, that we denote $c$ and $c_{t}$ respectively.
By Corollary 7.7, $(c_{t})_{t\in(1-\epsilon,1+\epsilon)}$ is a trivial deformation of $c=c_{1}$ and there exists a continuous family of isomorphisms $f_{t}:\mathrm{C}_{\mathcal{V}}\rightarrow\mathrm{C}_{\mathcal{V}_{t}}$ with $f_{1}=\mathrm{id}$. By fullness in Theorem 5.6, these isomorphisms lift to isomorphisms $\mathcal{V}\rightarrow\mathcal{V}_{t}$. Then, by the non-Abelian Hodge correspondence for complex modular/ribbon/braided functors (6.15), these isomorphisms induce isomorphisms $f_{t}:(\mathcal{V},\theta)\rightarrow(\mathcal{V},t\theta)$.
@@ 8. Proof of existence and uniqueness of Hodge structures on modular functors
By the correspondence (6.15) they induce automorphisms $k_{t}$ of $\mathcal{V}$ and thus also of the ribbon or braided fusion category $\mathrm{C}_{\mathcal{V}}$. By Corollary 7.10, there exists a family $\delta_{t}:\Lambda\rightarrow\mathbb{C}^{*}$ with $\delta_{t}(0)=0$ such that $k_{t}$ is on $\hom([\lambda],[\mu]\otimes[\nu])$ multiplication by $\delta_{t}(\mu)\delta_{t}(\nu)\delta_{t}(\lambda)^{-1}$. Thus $k_{t}$ is on $\mathcal{V}_{g}(\underline{\lambda})$ multiplication by $\prod_{i}\delta_{t}(\lambda_{i})$.
@@
8.2. Proof of Corollary 3.20
The proof essentially follows from that of [30, Theorem 5]. See there for details. Let $\mathcal{V}$ be a complex modular, genus 00 modular, ribbon or braided functor. Denote by $\Lambda$ its set of colors. Fix $g,n\geq 0$ tangent stable and $\underline{\lambda}\in\Lambda^{n}$. Denote by $\rho_{g}(\underline{\lambda})$ the monodromy representation of the local system $\mathcal{V}_{g}(\underline{\lambda})$, and by $\Gamma$ the fundamental group of ¯$\overline{\mathcal{M}}_{g}^{n}(r,s)$, ¯$\overline{\mathcal{M}}_{0}^{n}(r)$ or $^{1}\overline{\mathcal{M}}_{0,{n}}(r)$, depending on the type of functor considered.